22. Calculating the Cost of Debt Ying Import has several bond issues outstanding, each making semiannual interest payments. TCan you please show me how to do this with BA II Plus Financial Calculator? CCan you please show me how to do this with BA II Plus Financial Calculator?
Answer #1

Step 1:
N=5*2
PMT=6.5%*100/2
PV=-109
FV=100
CPT I/Y=2.24%

Step 2:
N=8*2
PMT=5.3%*100/2
PV=-97.4
FV=100
CPT I/Y=2.85%

Step 3:
N=15.5*2
PMT=6.9%*100/2
PV=-110.5
FV=100
CPT I/Y=2.93%

Step 4:
N=25*2
PMT=7.3%*100/2
PV=-109.8
FV=100
CPT I/Y=3.25%

Step 5:
=(30*109%*2.24%+50*97.4%*2.85%+65*110.5%*2.93%+85*109.8%*3.25%)/(30*109%+50*97.4%+65*110.5%+85*109.8%)*2*(1-24%)
=4.48%